OpenStudy (anonymous):

integral xcos(x^2)/square root of sin(x^2)

OpenStudy (anonymous):

Substitute sin(x^2) = t. Then 2xcos(x^2)dx = dt or, xcos(x^2)dx = dt/2 integral now becomes dt/(2*(square root of t)) solve using standard formulae now. Please click best response.